OpenStudy (anonymous):

you have a subscription to an online magazine that allows you to view 25 articles from the magazine's archives. you are charged an additional fee for each article after the first 25 articles viewed. after viewing 28 archived articles, you paid a total of $34.80. after viewing 30 archived articles, you paid a total of $40.70. a. what is the cost per archived article after the first 25 articles viewed? b. what is the cost of the magazine subscription?

OpenStudy (anonymous):

like finding the equation for the line through (3,34280) and (5,40.70) two more articles cost and additional 40.70-34.80 = 5.90 so the price per article is 5.90/2= $2.95 ( that is the slope)

OpenStudy (anonymous):

since you pay $2.95 per article after 25, and you were charged 34.80 for the three additional articles, you know you paid \[3\times 2.95 = 8.85\] for those three total, and 34.80 - 8.85 = 25.95 so the subscription cos $25.95 for 25 articles and an additional $2.95 for each article over 25